# Scaling Spoken Language Models with Syllabic Speech Tokenization

### v1

@@ 4 Results and Analysis
The results of our experiments are shown in Table 2. The table is broken into 3 parts, for each of the pretraining mixes that we decided to use. The first 3 columns are for the tokenizer, which show which SSL model it is based on, as well as the vocabulary size, and the total number of tokens in the training data. As we can see, the number of tokens overall is about 5x less for the Sylber based tokenizer compared to Hubert across the board. You can also see this in Table 1 which shows the statistics for the number of tokens for each of the datasets after tokenization. The 10 columns on the right are split into 2 sections, the first for the larger Qwen2.5-0.5B and the second for the smaller OPT-125M based model. Each section has 5 columns corresponding to the 5 different metrics detailed in Section 3.3.
@@ 4.1 Comparing Hubert and Sylber
The 5x reduction in context length can also be seen in the training time reduction. On an 8xA100-80GB NVIDIA DGX system, the final Hubert-based model trained on all 3 datasets takes 8.5 hours to complete while the Sylber KM2000 based model only takes 3 hours.

### v2

@@ 4 Results and Analysis
The results of our experiments are shown in Table 2. The table is broken into 3 parts, for each of the pretraining mixes that we decided to use. The first 3 columns are for the tokenizer, which show which SSL model it is based on, as well as the vocabulary size, and the total number of tokens in the training data. As we can see, the number of tokens overall is about 5x less for the Sylber based tokenizer compared to Hubert across the board. You can also see this in Table 1 which shows the statistics for the number of tokens for each of the datasets after tokenization. The 8 columns on the right are split into 2 sections, the first for the larger Qwen2.5-0.5B and the second for the smaller OPT-125M based model. Each section has 4 columns corresponding to the 4 different metrics detailed in Section 3.3.
@@ 4.1 Comparing Hubert and Sylber
The 5x reduction in context length can also be seen in the training time reduction. On an 8xA100-80GB NVIDIA DGX system, the final Hubert-based model trained on all 3 datasets takes 8.5 hours to complete while the Sylber KM20000 based model only takes 3 hours.